How many liters of a 25% alcohol solution must be mixed with a 12% solution to get 13L of a 15% solution?
Answer by richwmiller(17219)

a+b=13,
0.25*a+0.12*b=0.15*13
a=13-b
0.25*(13-b)+0.12*b=1.95
3.25-0.25b+0.12*b=1.95
-0.13*b=-1.3
b=10
a=13-b
a=3 liter at 25%
b=10 liter at 12%
check
0.25*3+0.12*10=0.15*13
0.75+1.2=1.95
1.95=1.95
ok
codemx1